# Estimating the Conformal Prediction Threshold from Noisy Labels

### v1

@@ Abstract
Conformal Prediction (CP) is a method to control prediction uncertainty by producing a small prediction set, ensuring a predetermined probability that the true class lies within this set. This is commonly done by defining a score, based on the model predictions, and setting a threshold on this score using a validation set. In this study, we address the problem of CP calibration when we only have access to a validation set with noisy labels. We show how we can estimate the noise-free conformal threshold based on the noisy labeled data. Our solution is flexible and can accommodate various modeling assumptions regarding the label contamination process, without needing any information about the underlying data distribution or the internal mechanisms of the machine learning classifier. We develop a coverage guarantee for uniform noise that is effective even in tasks with a large number of classes. We dub our approach Noise-Aware Conformal Prediction (NACP) and show on several natural and medical image classification datasets, including ImageNet, that it significantly outperforms current noisy label methods and achieves results comparable to those obtained with a clean validation set.
@@ 1 Introduction
In machine learning for safety-critical applications, the model must only make predictions it is confident about. One way to achieve this is by returning a (hopefully small) set of possible class candidates that contain the true class with a predefined level of certainty. This is a natural approach for medical imaging, where safety is of the utmost importance and a human makes the final decision. This allows us to aid the practitioner, by reducing the number of possible diagnoses he needs to consider, with a controlled chance of mistake. The general approach to return a prediction set without any assumptions on the data distribution (besides i.i.d. samples) is called Conformal Prediction (CP) [1, 19]. It creates a prediction set with the guarantee that the probability of the correct class being within this set meets or exceeds a specified confidence threshold. The goal is to return the smallest set possible while maintaining the confidence level guarantees. Recently, with the growing use of neural network systems in safety-critical applications such as medical imaging, CP has become an important calibration tool [11, 12, 14]. We note that CP is a general framework rather than a specific algorithm. The most common approach builds the prediction set using a conformity score, and different algorithms mostly vary in terms of how the conformity score is defined.
@@
When dealing with conformal predictions, a critical challenge arises in applications such as medical imaging due to label noise. In these domains, datasets frequently contain noisy labels stemming from ambiguous data that can confuse even clinical experts. Furthermore, physicians may disagree on the diagnosis for the same medical image, leading to inconsistencies in the ground truth labeling. Noisy labels also occur when applying differential privacy techniques to overcome privacy issues [5]. While significant efforts have been devoted to the problem of noise-robust network training [18, 21], the challenge of calibrating the models has only recently begun to receive attention.
@@
In this study, we tackle the challenge of applying CP to classification networks using a validation set with noisy labels. [3] suggested ignoring label noise and simply applying the standard CP algorithm on the noisy labeled validation set. This strategy results in large prediction sets especially when there are many classes. A recent study suggests estimating the noise-free conformal score given its noisy version and then applying the standard CP algorithm [15]. The most related study to ours is [17] which presents a noisy CP algorithm using conservative coverage guarantee bounds which can result in large prediction sets. Here, we present a novel algorithm for CP on noisy data that yields an effective coverage guarantee even in tasks with a large number of classes. We applied the algorithm to several standard medical and scenery imaging classification datasets and show that our method outperformed previous methods by a significant margin and achieved results comparable to those obtained by using a clean validation set.
@@ 2.1 Conformal Prediction
Consider a setup involving a classification network that categorizes an input $x$ into $k$ predetermined classes. Given a coverage level of $1-\alpha$, we aim to identify the smallest possible prediction set (a subset of these classes) ensuring the correct class is within the set with a probability of at least $1-\alpha$. A straightforward strategy to achieve this objective involves sequentially incorporating classes from the highest to the lowest probabilities until their cumulative sum exceeds the threshold of $1-\alpha$. Despite the network’s output adopting a mathematical distribution format, it does not inherently reflect the actual class distribution. Typically, the network will not be calibrated and it tends to be overly optimistic [6]. Consequently, this straightforward approach doesn’t assure the inclusion of the correct class with the desired probability.
The first step of the CP algorithm involves forming a conformity score $S(x,y)$ that measures the network’s uncertainty between $x$ and its true label $y$ (larger scores indicate worse agreement). The Homogeneous Prediction Sets (HPS) score [19] is $S_{\scriptscriptstyle\textrm{H}PS}(x,y)=1-p(y|x;\theta)$, s.t. $\theta$ is the network parameter set. The Adaptive Prediction Score (APS) [16] is the sum of all class probabilities that are not lower than the probability of the true class:
@@
such that $p_{i}=p(y=i|x;\theta)$ and $p_{y}$ is the probability of the label $y$. The RAPS score [2] is a variant of APS, which is defined as follows:
@@
$S_{\scriptscriptstyle rand-APS}(x,y,u)=\sum_{\{i|p_{i}>p_{y}\}}p_{i}+u\cdot p_{y},\hskip 28.45274ptu\sim U[0,1].$ (3)
The random version tends to yield the required coverage more precisely and thus it produces smaller prediction sets [1]. The CP prediction set of a data point $x$ is defined as $C_{q}(x)=\{y|S(x,y)\leq q\}$ where $q$ is a threshold that is found using a labeled validation set $(x_{1},y_{1}),...,(x_{n},y_{n})$. The CP theorem states that if we set $q$ to be the $(1\!-\!\alpha)$ quantile of the conformal scores $S(x_{1},y_{1}),...,S(x_{n},y_{n})$ we can guarantee that $1\!-\!\alpha\leq p(y\in C(x))\leq 1\!-\!\alpha+\frac{1}{n+1}$, where $x$ is a test point and $y$ is its the unknown true label [19]. In the random case there is still a coverage guarantee, which is defined by marginalizing over all test points $x$ and samplings $u$ from the uniform distribution [16]. Note that the coverage guarantee is for a marginal probability over all possible test points and coverage may be worse or better for different points. It can be proved that obtaining a conditional coverage guarantee is impossible [4].
3 Our Approach
@@ 3.1 Setting the Threshold Given Noisy Labels
Here we show how, given a simple noise model and a known noise level, we can get the correct CP threshold based on noisy data. We will generalize this beyond the simple noise model in the following section. Consider a network that classifies an input $x$ into $k$ pre-defined classes. Given a conformity score $S(x,y)$ and a specified coverage $1-\alpha$, the goal of the conformal prediction algorithm is to find a minimal $q$ such that $p(y\in C_{q}(x))\geq 1-\alpha$. Let $(x_{1},\tilde{y}_{1}),...,(x_{n},\tilde{y}_{n})$ be a validation set with noisy labels and let $y_{i}$ be the unknown correct label of $x_{i}$. Let $s_{i}=S(x_{i},\tilde{y}_{i})$ be the conformity score of $(x_{i},\tilde{y}_{i})$. We assume that the label noise follows a uniform distribution, where with a probability of $\epsilon$, the correct label is replaced by a label that is randomly sampled from the $k$ classes:
@@
Uniform noise is relevant, for example, when applying differential privacy techniques to overcome privacy issues [5]. In that setup the noise level $\epsilon$ is usually known. In other applications such as medical imaging, where the noise parameter $\epsilon$ is not given, it can be estimated with sufficient accuracy from the noisy-label data during training [24, 9, 10]. We can write $\tilde{y}$ as $\tilde{y}=(1-z)\cdot y+z\cdot u$, s.t. $u$ is a random label uniformly sampled from $\{1,...,k\}$ and $z$ is a binary random variable $(p(z=1)=\epsilon)$ indicating whether the label of the sample $(x,y)$ was replaced by a random label or not. For each candidate threshold $q$ denote:
@@
For each value $q$, we can estimate $F^{n}(q)$ from the noisy validation set:
@@
s.t. $u_{i}$ is uniformly sampled from $\{1,...,k\}$.
By substituting (6) and (7) in (5) we obtain an estimation of $F^{c}(q)=p(y\in C_{q}(x))$ based on the noisy validation set and the noise level $\epsilon$:
@@
For each candidate $q$ we first compute $\hat{F}^{n}(q)$ and $\hat{F}^{r}(q)$ and then by using (8) obtain the coverage estimation $\hat{F}^{c}(q)$. Given a coverage requirement $(1-\alpha)$, we can thus use the noisy validation set to find a threshold $q$ such that $\hat{F}^{c}(q)=1-\alpha$. Note that since $F^{c}(q)$ is monotonous, it seems reasonable to search for the threshold $q$ using the bisection method. However, as $\hat{F}^{c}(q)$ is an approximation based on the difference between two monotonic functions, it might not be exactly monotonous. We therefore find the threshold $q$ using an exhaustive grid search. We note that even with an exhaustive search the entire runtime is negligible compared to the training time. Furthermore, we can narrow the threshold search domain as follows:
@@
Finaly, $\hat{F}^{n}(q)$ is a monotonically increasing function of $q$ which implies that $q_{1}\leq q\leq q_{2}$. ∎
As an alternative to the grid search we can sort the noisy conformity scores $s_{i}=S(x_{i},\tilde{y}_{i})$ and look for the minimal $i$ such that $\hat{F}^{c}(s_{i})\geq 1-\alpha$. In the noise-free case $\hat{F}^{c}$ is piece-wise constant, with jumps determined exactly by the order statistics $s_{i}$, namely, $\hat{F}^{c}(s_{i})=i/n$ and thus this algorithm coincides with the standard CP algorithm. In the noisy case $\hat{F}^{c}(q)$ depends on the conformity scores of all the $k$ classes and thus its structure is more complicated. We dub our algorithm Noise-Aware Conformal Prediction (NACP), and summarize it in Algorithm Box 1. Note that in the noise-free case ($\epsilon=0$) the NACP algorithm coincides with the standard CP algorithm and selects $q$ that satisfies $\hat{F}^{c}(q)=\hat{F}^{n}(q)=1-\alpha$, i.e., $q$ is the $1-\alpha$ quantile of the validation set conformity scores.
@@ 3.2 Prediction Size Comparison
We next compare our NACP approach analytically to Noisy-CP [3] in terms of the average size of the prediction set.
@@
Hence from the monotonicity of $\hat{F}^{n}(q)$ we have $q\leq\tilde{q}$ iff $\hat{F}^{n}(q)\leq\hat{F}^{n}(\tilde{q})$ iff $\hat{F}^{r}({q})\leq 1-\alpha$. ∎
The theorem above states that if the size of the prediction set obtained by NACP is less than $k(1-\alpha)$, NACP is more effective than Noisy-CP. For example, assume $k=100$ and $0.91-\alpha=$ 0.9. In this case, if the average size of the NACP prediction set is less than 90, NACP is more effective than Noisy-CP. We also see from eq. (11) that the smaller $\hat{F}^{r}$ is the larger the gap between the two methods. Since $\hat{F}^{r}$ is inversely proportional to the number of classes, we expect the difference to be substantial when there is a large number of classes to consider, which is exactly where CPs’ ability to reliably exclude possible classes is very useful. In our experiments, we indeed found a considerable gap between the two methods when we experimented on classification tasks with a large number of classes.
@@
1:Input: A conformity score $S(x,y)$, a coverage level $1\!-\!\alpha$ and a validation set $(x_{1},\tilde{y}_{1}),...,(x_{n},\tilde{y}_{n})$, s.t. the labels are corrupted by a uniform noise with parameter $\epsilon$.
@@
$\hat{F}^{n}(q)=\frac{1}{n}\sum_{i}\mathds{1}_{\{\tilde{y}_{i}\in C_{q}(x_{i})\}},\hskip 19.91684pt\hat{F}^{r}(q)=\frac{1}{n}\sum_{i}\frac{|C_{q}(x_{i})|}{k},\hskip 19.91684pt\hat{F}^{c}(q)=\frac{\hat{F}^{n}(q)-\epsilon\hat{F}^{r}(q)}{1-\epsilon}$
@@
5:The prediction set of a test sample $x$ is $C_{q}(x)=\{y\,|\,S(x,y)<q\}$.
6:Coverage guarantee: $p(y\in C_{q}(x))\geq 1-\alpha-\Delta(n,\epsilon,\delta)$ with probability $(1-\delta)$ over the noisy validation set sampling (see Theorem 3.5).
@@ 3.3 Coverage Guarantees
We next provide a coverage guarantee for NACP. We show that if we apply the NACP to find a threshold $q$ for $1-\alpha+\Delta$, then $P(y\in C_{q}(x))\geq 1-\alpha$ were $\Delta$ depends on the validation set size. $\Delta$ is a finite-sample term that is needed to approximate the CDF to set the threshold instead of simply picking a predefined quantile. Because $\Delta$ can be computed, one can adjust the $\alpha$ used in the NACP algorithm to get the desired coverage guarantee. However, we note that we empirically found this bound to be over-conservative, and that the un-adjusted method does reach the desired coverage.
@@
Given $\delta>0$, define $\Delta=\sqrt{\frac{\log(4/\delta)}{2nh^{2}}}$ such that $h=\frac{1-\epsilon}{1+\epsilon}$ and $n$ is the size of the noisy validation set. Then
@@
such that the probability is over the validation set.
Proof.
The Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [13] states that if we estimate a CDF $F$ from $n$ samples using the empirical CDF $F_{n}$ then $p(\sup_{x}|F_{n}(x)-F(x)|>\Delta)\leq 2\exp(-2n\Delta^{2}$). Eq. (8) defines $\hat{F}^{c}(q)$ using $\hat{F}^{n}(q)$ and $\hat{F}^{r}(q)$. Both are empirical CDF, so from the DKW theorem and the union bound we get that:
$p(\sup_{q}|F^{r}(q)-\hat{F}^{r}(q)|>h\Delta\,\,\mbox{or}\,\,\sup_{q}|F^{n}(q)-\hat{F}^{n}(q)|>h\Delta)\leq 4\exp(-2nh^{2}\Delta^{2})=\delta.$ (13)
@@
$\hat{F}^{c}(q)=\frac{\hat{F}^{n}(q)-\epsilon\hat{F}^{r}(q)}{1-\epsilon}\leq\frac{(F^{n}(q)+h\Delta)-\epsilon(F^{r}(q)-h\Delta)}{1-\epsilon}$ (14)
$=F^{c}(q)+\frac{h\Delta+\epsilon h\Delta}{1-\epsilon}=F^{c}(q)+h\Delta\frac{1+\epsilon}{1-\epsilon}=F^{c}(q)+\Delta.$
Similarly, we can show that $\hat{F}^{c}(q)\geq F^{c}(q)-\Delta$ which completes the proof. ∎
@@
Assume you have a noisy validation set of size $n$ with noise level $\epsilon$ and set $\Delta(n,\epsilon,\delta)=\sqrt{\frac{\log(4/\delta)}{2nh^{2}}}$ s.t. $h=\frac{1-\epsilon}{1+\epsilon}$ and that you pick $q$ such that $\hat{F}^{c}(q)=1-\alpha+\Delta$. Then with probability at least $1-\delta$ (over the validation set), we have that if $(x,y)$ are sampled from the clear label distribution we get:
$1-\alpha\leq p(y\in C_{q}(x))\leq 1-\alpha+2\Delta.$
Proof.
Given a clean test pair $(x,y)$, with probability $\delta$ over the validation set, we have:
$p(y\in C_{q}(x))=p(S(x,y)<q)=F^{c}(q)\geq\hat{F}^{c}(q)-\Delta=1-\alpha.$
In a similar way: $p(y\in C_{q}(x))=F^{c}(q)\leq\hat{F}^{c}(q)+\Delta=1-\alpha+2\Delta$. ∎
As the size of the noisy validation set, $n$, tends to infinity, $\Delta$ converges to zero and thus the noisy threshold converges to the noise-free threshold.
Algorithm 2 Noise-Aware Conformal Prediction (NACP) for a noise matrix model
1:Input: A conformity score $S(x,y)$, a coverage level $1\!-\!\alpha$ and a validation set $(x_{1},\tilde{y}_{1}),...,(x_{n},\tilde{y}_{n})$, s.t. the labels are corrupted by a noise matrix $P$.
2:For each candidate threshold $q$ compute:
$\hat{M}_{q}(\ell,i)=\frac{1}{n}\sum_{j}\mathds{1}_{\{\tilde{y}_{j}=i,\,\,\ell\in C_{q}(x_{j})\}},\hskip 14.22636pti,\ell=1,..,k.$
$\hat{F}^{c}(q)=\textrm{Tr}(\hat{M}_{q}P^{-1}).$
3:Apply a grid search to find $q$ that satisfies $\hat{F}^{c}(q)=1\!-\!\alpha$.
4:The prediction set of a test sample $x$ is $C_{q}(x)=\{y\,|\,S(x,y)<q\}$.
[17] proposed a CP algorithm for the same setup of noisy labels where the noise matrix is given (or estimated based on noise-free data). They used the distribution of correct labels given the noisy labels while we use the more natural distribution of the noisy labels given the correct label. As a result, they need to know the class frequencies for both the clean and noisy labels, whereas we do not. Another major difference between the algorithms is the finite sample coverage guarantee term. In Section 4 we empirically compare the finite sample terms of the two methods and show that while ours is effective for the case of uniform noise, their approach fails on datasets with a large number of classes such as CIFAR-100, TinyImageNet, and ImageNet, and produces prediction sets that contain all classes. A further distinction is that their finite sample coverage guarantee is established for the average of all the noisy validation sets. In contrast, our approach provides an individual coverage guarantee for nearly all ($1-\delta$) of the sampled noisy validation sets.

### v2

# Conformal Prediction of Classifiers with Many Classes based on Noisy Labels

@@ Abstract
Conformal Prediction (CP) controls the prediction uncertainty of classification systems by producing a small prediction set, ensuring a predetermined probability that the true class lies within this set. This is commonly done by defining a score, based on the model predictions, and setting a threshold on this score using a validation set. In this study, we address the problem of CP calibration when we only have access to a calibration set with noisy labels. We show how we can estimate the noise-free conformal threshold based on the noisy labeled data. We derive a finite sample coverage guarantee for uniform noise that remains effective even in tasks with a large number of classes. We dub our approach Noise-Aware Conformal Prediction (NACP). We illustrate the performance of the proposed results on several standard image classification datasets with a large number of classes.
@@ 1 Introduction
In machine learning for safety-critical applications, the model must only make predictions it is confident about. One way to achieve this is by returning a (hopefully small) set of possible class candidates that contain the true class with a predefined level of certainty. This is a natural approach for medical imaging, where safety is of the utmost importance and a human makes the final decision. This allows us to aid the practitioner, by reducing the number of possible diagnoses he needs to consider, with a controlled risk of mistake. The general approach to return a prediction set without any assumptions on the data distribution (besides i.i.d. samples) is called Conformal Prediction (CP) (Angelopoulos et al., 2023; Vovk et al., 2005). It creates a prediction set with the guarantee that the probability of the correct class being within this set meets or exceeds a specified confidence threshold. The goal is to return the smallest set possible while maintaining the confidence level guarantees. Recently, with the growing use of neural network systems in safety-critical applications such as medical imaging, CP has become an important calibration tool (Lu et al., 2022a, b; Olsson et al., 2022). We note that CP is a general framework rather than a specific algorithm. The most common approach builds the prediction set using a conformity score, and different algorithms mostly vary in terms of how the conformity score is defined.
@@
When dealing with conformal predictions, a critical challenge arises in applications such as medical imaging due to label noise. In these domains, datasets frequently contain noisy labels stemming from ambiguous data that can confuse even clinical experts. Furthermore, physicians may disagree on the diagnosis for the same medical image, leading to inconsistencies in the ground truth labeling. Noisy labels also occur when applying differential privacy techniques to overcome privacy issues (Ghazi et al., 2021; Penso et al., 2025). While significant efforts have been devoted to the problem of noise-robust network training (Song et al., 2022; Xue et al., 2022), the challenge of calibrating the models has only recently begun to receive attention (Penso et al., 2024).
@@
In this study, we tackle the challenge of applying CP to classification networks using a calibration set with noisy labels. Einbinder et al. (2022) suggested ignoring label noise and simply applying the standard CP algorithm on the noisy labeled calibration set. This strategy results in large prediction sets especially when there are many classes. A recent study suggests estimating the noise-free conformal score given its noisy version and then applying the standard CP algorithm (Penso and Goldberger, 2024). The most related studies to ours are (Sesia et al., 2024) and (Clarkson et al., 2024) which present a noisy CP algorithm with coverage guarantee bounds that can be too conservative in tasks with many classes (a detailed discussion on related works appears in Section 4). Here, we present a novel algorithm for CP on noisy data that yields an effective coverage guarantee even in tasks with a large number of classes. We applied the algorithm to several standard medical and scenery imaging classification datasets and show that our method outperformed previous methods.
@@ 2.1 Conformal Prediction
Consider a setup involving a classification network that categorizes an input $x$ into $k$ predetermined classes. Given a coverage level of $1-\alpha$, we aim to identify the smallest possible prediction set (a subset of these classes) ensuring the correct class is within the set with a probability of at least $1-\alpha$. A straightforward strategy to achieve this objective involves sequentially incorporating classes from the highest to the lowest probabilities until their cumulative sum exceeds the threshold of $1-\alpha$. Despite the network’s output adopting a mathematical distribution format, it does not inherently reflect the actual class distribution. Typically, the network will not be calibrated and it tends to be overly optimistic (Guo et al., 2017). Consequently, this straightforward approach doesn’t assure the inclusion of the correct class with the desired probability.
The first step of the CP algorithm involves forming a conformity score $S(x,y)$ that measures the network’s uncertainty between $x$ and its true label $y$ (larger scores indicate worse agreement). The Homogeneous Prediction Sets (HPS) score (Vovk et al., 2005) is $S_{\scriptscriptstyle\textrm{H}PS}(x,y)=1-p(y|x;\theta)$, s.t. $\theta$ is the network parameter set. The Adaptive Prediction Score (APS) (Romano et al., 2020) is the sum of all class probabilities that are not lower than the probability of the true class:
@@
such that $p_{i}=p(y=i|x;\theta)$ and $p_{y}$ is the probability of the label $y$. The RAPS score (Angelopoulos et al., 2021) is a variant of APS, which is defined as follows:
@@
$S_{\scriptscriptstyle APS}(x,y,u)=\sum_{\{i|p_{i}>p_{y}\}}p_{i}+u\cdot p_{y},\hskip 8.5359ptu\sim U[0,1].$ (3)
The random version tends to yield the required coverage more precisely and thus it produces smaller prediction sets (Angelopoulos et al., 2023). The CP prediction set of a data point $x$ is defined as $C_{q}(x)=\{y|S(x,y)\leq q\}$ where $q$ is a threshold that is found using a labeled calibration set $(x_{1},y_{1}),...,(x_{n},y_{n})$. The CP theorem (Vovk et al., 2005) states that if we set $q$ to be the $(1\!-\!\alpha)$ quantile of the conformal scores $S(x_{1},y_{1}),...,S(x_{n},y_{n})$ we can guarantee that $1\!-\!\alpha\leq p(y\in C(x))$. where $x$ is a test point and $y$ is its the unknown true label (Vovk et al., 2005). In the random case there is still a coverage guarantee, which is defined by marginalizing over all test points $x$ and samplings $u$ from the uniform distribution (Romano et al., 2020).
3 CP Calibration based on Noisy Labels
@@ 3.1 Setting the Threshold Given Noisy Labels
Here we show how, given a simple noise model and a known noise level, we can get the correct CP threshold based on noisy data. We will generalize this beyond the simple noise model in the following section. Consider a network that classifies an input $x$ into $k$ pre-defined classes. Given a conformity score $S(x,y)$ and a specified coverage $1-\alpha$, the goal of the conformal prediction algorithm is to find a minimal $q$ such that $p(y\in C_{q}(x))\geq 1-\alpha$. Let $(x_{1},\tilde{y}_{1}),...,(x_{n},\tilde{y}_{n})$ be a calibration set with noisy labels and let $y_{i}$ be the unknown correct label of $x_{i}$. Let $s_{i}=S(x_{i},\tilde{y}_{i})$ be the conformity score of $(x_{i},\tilde{y}_{i})$. We assume that the label noise follows a uniform distribution, where with a probability of $\epsilon$, the correct label is replaced by a label that is randomly sampled from the $k$ classes:
@@
Uniform noise is relevant, for example, when applying differential privacy techniques to overcome privacy issues (Ghazi et al., 2021). In that setup the noise level $\epsilon$ is usually known. In other applications such as medical imaging, where the noise parameter $\epsilon$ is not given, it can be estimated with sufficient accuracy from the noisy-label data during training (Zhang et al., 2021; Li et al., 2021; Lin et al., 2023). We can write $\tilde{y}$ as $\tilde{y}=(1-z)\cdot y+z\cdot u$, where $u$ is a random label uniformly sampled from $\{1,...,k\}$ and $z$ is a binary random variable $(p(z=1)=\epsilon)$ indicating whether the label of the sample $(x,y)$ was replaced by a random label or not. For each candidate threshold, $q$ denote:
@@
For each value $q$, we can estimate $F^{n}(q)$ from the noisy calibration set:
@@
where $u_{i}$ is uniformly sampled from $\{1,...,k\}$. By substituting (6) and (7) in (5) we obtain an estimation of $F^{c}(q)=p(y\in C_{q}(x))$ based on the noisy calibration set and the noise level $\epsilon$:
@@
For each candidate $q$ we first compute $\hat{F}^{n}(q)$ and $\hat{F}^{r}(q)$ and then by using (8) obtain the coverage estimation $\hat{F}^{c}(q)$. Given a coverage requirement $(1-\alpha)$, we can thus use the noisy calibration set to find a threshold $q$ such that $\hat{F}^{c}(q)=1-\alpha$. Note that since $F^{c}(q)$ is monotonous, it seems reasonable to search for the threshold $q$ using the bisection method. However, as $\hat{F}^{c}(q)$ is an approximation based on the difference between two monotonic functions, it might not be exactly monotonous. We therefore find the threshold $q$ using an exhaustive grid search. If there are several solutions we select the largest value. (In practice selecting one of the solutions has almost no effect on the results.) We note that even with an exhaustive search the entire runtime is negligible compared to the training time. We can narrow the threshold search domain as follows:
@@
Finaly, $\hat{F}^{n}(q)$ is a monotonically increasing function of $q$ which implies that $q_{1}\leq q\leq q_{2}$.
As an alternative to the grid search we can sort the noisy conformity scores $s_{i}=S(x_{i},\tilde{y}_{i})$ and look for the minimal $i$ such that $\hat{F}^{c}(s_{i})\geq 1-\alpha$. In the noise-free case $\hat{F}^{c}$ is piece-wise constant, with jumps determined exactly by the order statistics $s_{i}$, namely, $\hat{F}^{c}(s_{i})=i/n$ and thus this algorithm coincides with the standard CP algorithm. We dub our algorithm Noise-Aware Conformal Prediction (NACP), and summarize it in Algorithm 1. Note that in the noise-free case ($\epsilon=0$) the NACP algorithm coincides with the standard CP algorithm and selects $q$ that satisfies $\hat{F}^{c}(q)=\hat{F}^{n}(q)=1-\alpha$, i.e., $q$ is the $1-\alpha$ quantile of the calibration set conformity scores.
@@ 3.2 Prediction Size Comparison
We next compare our NACP approach analytically to Noisy-CP (Einbinder et al., 2022) in terms of the average size of the prediction set.
@@
[…]
The theorem above states that if the size of the prediction set obtained by NACP is less than $k(1-\alpha)$, NACP is more effective than Noisy-CP. For example, assume $k=100$ and $1-\alpha=0.9$. In this case, if the average size of the NACP prediction set is less than 90, NACP is more effective than Noisy-CP. We also see from eq. (11) that the smaller $\hat{F}^{r}$ is the larger the gap between the two methods. Since $\hat{F}^{r}$ is inversely proportional to the number of classes, we expect the difference to be substantial when there is a large number of classes to consider, which is exactly where CPs’ ability to reliably exclude possible classes is very useful.
@@
1:Input: A conformity score $S(x,y)$, a coverage level $1\!-\!\alpha$ and a calibration set $(x_{1},\tilde{y}_{1}),...,(x_{n},\tilde{y}_{n})$, where the labels are corrupted by a uniform noise with parameter $\epsilon$.
@@
$\hat{F}^{n}(q)=\frac{1}{n}\sum_{i}\mathds{1}_{\{\tilde{y}_{i}\in C_{q}(x_{i})\}},\hskip 14.22636pt\hat{F}^{r}(q)=\frac{1}{n}\sum_{i}\frac{|C_{q}(x_{i})|}{k},\hskip 14.22636pt\hat{F}^{c}(q)=\frac{\hat{F}^{n}(q)-\epsilon\hat{F}^{r}(q)}{1-\epsilon}$
@@
5:The prediction set of a test sample $x$ is:
$C_{q}(x)=\{y\,|\,S(x,y)\leq q\}.$
6:Coverage guarantee: $p(y\in C_{q}(x))\geq 1-\alpha-\Delta(n,\epsilon,\delta)$ with probability $(1-\delta)$ over the noisy calibration set sampling (see Theorem 3.9).
@@ 3.3 Coverage Guarantees
We next provide a coverage guarantee for NACP. We show that if we apply the NACP to find a threshold $q$ for $1-\alpha+\Delta$, then $P(y\in C_{q}(x))\geq 1-\alpha$ were $\Delta$ depends on the calibration set size. $\Delta$ is a finite-sample term that is needed to approximate the CDF to set the threshold instead of simply picking a predefined quantile. Because $\Delta$ can be computed, one can adjust the $\alpha$ used in the NACP algorithm to get the desired coverage guarantee. We note that we empirically found this bound to be over-conservative and that the un-adjusted method does reach the desired coverage (see Section 5).
@@
Given $\delta>0$, define $\Delta=\sqrt{\frac{\log(4/\delta)}{2nh^{2}}}$ such that $h=\frac{1-\epsilon}{1+\epsilon}$ and $n$ is the size of the noisy calibration set. Then
@@
such that the probability is over the sampling of the noisy calibration set.
Proof 3.8.
The Dvoretzky–Kiefer–Wolfowitz (DKW) inequality (Massart, 1990) states that if we estimate a CDF $F$ from $n$ samples using the empirical CDF $F_{n}$ then $p(\sup_{x}|F_{n}(x)-F(x)|>\Delta)\leq 2\exp(-2n\Delta^{2}$). Eq. (8) defines $\hat{F}^{c}(q)$ using $\hat{F}^{n}(q)$ and $\hat{F}^{r}(q)$. Both are empirical CDF, so from the DKW theorem and the union bound we get that:
$p(\sup_{q}|F^{r}(q)-\hat{F}^{r}(q)|>h\Delta\,\,\mbox{or}\,\,\sup_{q}|F^{n}(q)-\hat{F}^{n}(q)|>h\Delta)\leq 4\exp(-2nh^{2}\Delta^{2})=\delta.$
@@
$\hat{F}^{c}(q)=\frac{\hat{F}^{n}(q)-\epsilon\hat{F}^{r}(q)}{1-\epsilon}\leq\frac{(F^{n}(q)+h\Delta)-\epsilon(F^{r}(q)-h\Delta)}{1-\epsilon}F^{c}(q)+\frac{h\Delta+\epsilon h\Delta}{1-\epsilon}$
$=F^{c}(q)+h\Delta\frac{1+\epsilon}{1-\epsilon}=F^{c}(q)+\Delta.$
Similarly, we can show that $\hat{F}^{c}(q)\geq F^{c}(q)-\Delta$ which completes the proof.
@@
Assume you have a noisy calibration set of size $n$ with noise level $\epsilon$ and set $\Delta(n,\epsilon,\delta)=\sqrt{\frac{\log(4/\delta)}{2nh^{2}}}$ where $h=\frac{1-\epsilon}{1+\epsilon}$ and that you pick $q$ such that $\hat{F}^{c}(q)\geq 1-\alpha+\Delta$. Then with probability at least $1-\delta$ (over the calibration set), we have that if $(x,y)$ are sampled from the clear label distribution we get:
$1-\alpha\leq p(y\in C_{q}(x)).$
Proof 3.10.
Given a clean test pair $(x,y)$, with probability $\delta$ over the calibration set, we have:
$p(y\in C_{q}(x))=p(S(x,y)<q)=F^{c}(q)\geq\hat{F}^{c}(q)-\Delta\geq 1-\alpha.$
As the size of the noisy calibration set, $n$, tends to infinity, $\Delta$ converges to zero and thus the noisy threshold converges to the noise-free threshold.